OpenStudy (anonymous):

Lisa is driving a car at an average speed of 55 mi/h. A. What is Lisa's average speed in feet per second. B. How many feet will Lisa travel in 40 min? Show your work

OpenStudy (anonymous):

Wow thank you.!

OpenStudy (anonymous):

wait were did it go

OpenStudy (texaschic101):

55 miles per hr 5280 ft per mile 3600 seconds per hr (5280(55)) / 3600 = 290400/3600 = 80.67 ft per second (A) =============== travels 80.67 ft per second 40 minutes = 2400 seconds 80.67/1 = x/2400 -- (80.67 ft to 1 sec = x ft to 2400 seconds) cross multiply x = 2400(80.67) x = 193608 ft in 40 minutes (B) I think I did the second part right...you might need to get a second opinion